OpenStudy (anonymous):

giving best answer or fanning solve x+3/(3x)>2

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

\[x+3>6x\]

hartnn (hartnn):

not actually, how do you know whether x is positive or negative, and whether to switch signs or not ?

hartnn (hartnn):

the question is \(\dfrac{x+3}{3x}>2\) right ?

OpenStudy (anonymous):

cant we write x>-3/5?

OpenStudy (anonymous):

yes that is the question @hartnn

hartnn (hartnn):

then the only way to do that is to break up, \(\dfrac{x}{3x}+\dfrac{3}{3x}>2\)

hartnn (hartnn):

then 'x' gets cancelled, and you can move 1/3 to right side because its a constant, not a variable

hartnn (hartnn):

@Alexis2332 try to contibue ?

hartnn (hartnn):

**continue

OpenStudy (anonymous):

not 100% what to do next

hartnn (hartnn):

\(\dfrac{1}{3}+\dfrac{3}{3x}>2\) subtract 1/3 from both sides, what u get ?

OpenStudy (anonymous):

3/3x > 1 1/3

hartnn (hartnn):

2-1/3 = 1 1/3 or 1 2/3 ??

hartnn (hartnn):

2-1/3 = (2*3-1 ) /3 = 5/3 = (3+2)/3 = 1 2/3 got this ?

OpenStudy (anonymous):

yes got that

hartnn (hartnn):

ok, also 3/3x simplifies to 1/3 (the 3's cancel out) and lets keep that as 5/3 , so we have \(\dfrac{1}{x}>\dfrac{5}{3}\) now just take the reciprocal, and flip the sign! like if \(\dfrac{a}{b}>\dfrac{c}{d} \implies \dfrac{b}{a}>\dfrac{d}{c} \) so ,what about \(\dfrac{1}{x}>\dfrac{5}{3} \) ???

hartnn (hartnn):

**** \(\dfrac{a}{b}>\dfrac{c}{d} \implies \dfrac{b}{a}<\dfrac{d}{c}\)

OpenStudy (anonymous):

x/1 < 3/5

hartnn (hartnn):

x < 3/5 correct

hartnn (hartnn):

and to make it more accurate, \(0<x<3/5\)

OpenStudy (anonymous):

oh okay I got it, thanks!

hartnn (hartnn):

welcome ^_^